OpenStudy (anonymous):

The temperature of magnesium must be at least 650 degrees for the metal to melt. A sample of magnesium is currently 255 degrees. The temperature of the magnesium is increased by 10 degrees every minute. Which inequality shows the number of minutes (m) needed to raise the temperature to the melting point? m > 39.5 m ≥ 39.5 m > 87.5 m ≥ 87.5

OpenStudy (texaschic101):

255 + 10m >= 650 10m >= 650 - 255 10m >= 395 m >= 395/10 m >= 39.5